40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 7 09:00:26 2022, Paco, Update, General, Scheduled power outage recovery 
    Reply  Mon Feb 7 15:32:48 2022, Yehonathan, Update, General, Y End laser controller 
    Reply  Mon Feb 7 16:53:02 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220208_001646282.jpgPXL_20220208_001657871.jpg
       Reply  Wed Feb 9 11:56:24 2022, Anchal, Update, General, Bringing back CDS Screenshot_2022-02-09_12-11-33.png
          Reply  Wed Feb 9 13:55:05 2022, Koji, Update, General, Bringing back CDS 
       Reply  Wed Feb 9 16:43:35 2022, Paco, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
          Reply  Thu Feb 10 15:41:00 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
             Reply  Thu Feb 10 17:57:48 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) PXL_20220211_021509819.jpg
                Reply  Thu Feb 10 19:03:23 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                   Reply  Thu Feb 10 19:46:37 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                      Reply  Thu Feb 10 21:10:43 2022, Koji, Update, General, Video Mux setting reset Screenshot_2022-02-10_21-11-21.pngScreenshot_2022-02-10_21-11-54.png
             Reply  Fri Feb 11 16:09:11 2022, Anchal, Update, General, Scheduled power outage recovery - Input power increased 
                Reply  Mon Feb 14 18:31:50 2022, Paco, Update, General, Scheduled power outage recovery - IMC recovery progress 
                   Reply  Tue Feb 15 19:32:50 2022, Koji, Update, General, Scheduled power outage recovery - IMC recovery progress PXL_20220216_001731377.jpgScreen_Shot_2022-02-15_at_16.18.16.pngPXL_20220216_001727465.jpgScreen_Shot_2022-02-15_at_16.22.16.pngPXL_20220216_002229572.jpg
                      Reply  Tue Feb 15 19:40:02 2022, Koji, Update, General, IMC locking 
                         Reply  Wed Feb 16 15:19:41 2022, Anchal, Update, General, Reconfigured MC reflection path for low power 
                            Reply  Wed Feb 23 15:08:57 2022, Anchal, Update, General, Removed extra beamsplitter in MC WFS path 
                               Reply  Thu Feb 24 14:32:57 2022, Anchal, Update, General, MC RFPD DCMON channel got stuck to 0 
                                  Reply  Thu Feb 24 19:26:32 2022, Anchal, Update, General, IMC Locking 
                                     Reply  Sun Feb 27 00:37:00 2022, Koji, Update, General, IMC Locking Recovery PXL_20220226_093809056.jpgPXL_20220226_093854857.jpgPXL_20220226_100859871.jpgScreenshot_2022-02-26_01-56-31.pngScreenshot_2022-02-26_01-56-47.png
                                        Reply  Sun Feb 27 01:12:46 2022, Koji, Update, General, IMC manual alignment procedure PXL_20220226_100859871.jpg
    Reply  Fri Feb 11 11:17:00 2022, Anchal, Update, General, Scheduled power outage recovery 
       Reply  Mon Feb 14 21:03:25 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220215_025325118.jpg
Message ID: 16648     Entry time: Mon Feb 7 09:00:26 2022     Reply to this: 16649   16651   16665
Author: Paco 
Type: Update 
Category: General 
Subject: Scheduled power outage recovery 

[Paco]

Started recovering from scheduled (Feb 05) power outage. Basically, time-reversing through this list.


== Office area ==

  • Power martian network switches, WiFi routers on the north-rack.
  • Power windows (CAD) machine on.

== Main network stations ==

  • Power on nodus, try ping (fail).
  • Power on network switches, try ping (success), try ssh controls@nodus.ligo.caltech.edu (success).
  • Power on chiara to serve names for other stations, try ssh chiara (success).
  • Power on fb1, try ping (success), try ssh fb1 (success).
  • Power on paola (xend laptop), viviana (yend laptop), optimus, megatron.

== Control workstations ==

  • Power on zita (success)
  • Power on giada (success), run system upgrade.
  • Power on donatella (success)
  • Power on allegra (fail)  **
  • Power on pianosa (success)
  • Power on rossa (success)
  • From nodus, started elog (success).

== PSL + Vertex instruments ==

  • Turn on newport PD power supplies on PSL table.
  • Turn on TC200 temp controller on (setpoint --> 36.9 C)
  • Turn on two oscilloscopes in PSL table.
  • Turn on PSL (current setpoint --> 2.1 A, other settings seem nominal)
  • Turn on Thorlabs HV pzt supply.
  • Turn on ITMX OpLev / laser instrument AC strip.

== YEND and XEND instruments ==

  • Turn on XEND AUX pump on (current setpoint -->1.984 A)
  • Turn on XEND AUX SHG oven on (setpoint --> 37.1 C) (see green beam)
  • Turn on XEND AUX shutter controller on.
  • Turn on DCPD supply, and OpLev supply AC strip on.
  • Turn on YEND AUX pump on (fail) *
    • With the controller on STDBY, I tried setting up the current but got HD FAULT (or according to the manual this is what the head reports when the diode temperature is too high...)
    • Upon power cycling the controller, even the controller display stopped working... YAUX controller + head died? maybe just the diode? maybe just the controller?
      • I borrowed a spare LW125 controller from the PSL table (Yehonathan pointed me to it) and swapped it in.
      • Got YEND AUX to lase with this controller, so the old controller is busted but at least the laser head is fine.
      • Even saw SHG light. We switched the laser head off to "STDBY" (so it remains warm) and took the faulty controller out of there.
  • Turn on YEND AUX SHG oven on (setpoint -->35.7 C)
  • Turn on YEND AUX shutter controller on.

== YARM Electronic racks ==

== XARM Electronic racks ==

 


* Top priority, this needs to be fixed.

** Non-priority, but to be debugged

ELOG V3.1.3-